Solve using elimination process.
x+2y=6
2x-3y=26

Answer by pie90010(44)

x+2y=6
2x-3y=26
so you want to eliminate the Xs , its a lot easier
to do this, just multiply the whould first equation by -2
-2x - 4y = -12
2x - 3y = 26
X's cancel, now just add it upp
-7y = 14
y = -7
then to get X, plug it into the original equation
x + 2(-7) = 6
x - 14 = 6
x = 20
the salution would be (20, -7)